If the sum of the ages of two students is 40, and one is two years older then there other, what are both of their ages? Show in an equation, not the answer.
Answer by dkppathak(439)

If the sum of the ages of two students is 40, and one is two years older then there other, what are both of their ages? Show in an equation, not the answer.
solution
let once age is x yrs and others 2 yrs older means other age will be x+2
as per given condition sum of both age is 40 yrs means
x+x+2=40
2x+2=40 this is equation or 2(x+1)=40 or x+1=20 or x=20-1=19
so one is 19 yrs and other is 21 yrs